OpenStudy (anonymous):

HELP! A rectangular athletic field is twice a long as it is wide. If the perimeter of the athletic field is 216 yards, what are its dimensions?

OpenStudy (aravindg):

area of field is =2(l+b) where l is length b is breadth here l=2b so (2(2b+b)=216 6b=216 b=36

OpenStudy (anonymous):

I'm still a bit confused...I tried using the equation 2l+2w

OpenStudy (aravindg):

its the same 2l+2w =2(l+w) only diff is i called width as b you call it as w

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

Yeah I think I got it. Thank you!